OpenStudy (anonymous):

Area problem:

OpenStudy (zzr0ck3r):

find h with Pythagorean theorem

OpenStudy (anonymous):

427.4 is what I got.

OpenStudy (zzr0ck3r):

\(SA=\pi rs+\pi r^2\) and we remove the circle top so \(SA=\pi r s\)

OpenStudy (zzr0ck3r):

correct